OpenStudy (anonymous):

Which polynomial is prime? A. 4x^2 - 17x + 15 B. 4x^2 + 13x C. 8x^2 + 7x - 16 D. x^2 - 4

jhonyy9 (jhonyy9):

(x-2)(x+2)

OpenStudy (turingtest):

Hi meow!

OpenStudy (anonymous):

hi !!!!!

OpenStudy (anonymous):

So I'm really confused on this.. how does (x-2)(x+2) come into the picture for this equation?

OpenStudy (turingtest):

jhonyy factored the last expression x^2 - 4=(x+2)(x-2) because we can always factor a^2-b^2=(a-b)(a+b) that is called difference of squares. The answer to you question, however, is C, because that polynomial is not factorable, so it is prime.

OpenStudy (anonymous):

Oh! Alright because I was just super confused at first and I'm like... hm... lol. How was your New Year's?

OpenStudy (turingtest):

jolly, yours?

OpenStudy (anonymous):

It was great! Got my hair re-done over the weekend!

OpenStudy (turingtest):

Looks great! well ,here's the idea of what you are doing factor each answer: 4x^2 - 17x + 15=(x-3)(4x-5) 4x^2 + 13x=x(4x+13) 8x^2 + 7x - 16=NOT FACTORABLE (not traditionally at least) x^2 - 4=(x+2)(x-2) since the third one is the only one we can't factor at all, the answer is C See ya 'round!

OpenStudy (anonymous):

Thank you! (: And thanks for the help! (: You are awesome, just saying! See you around too! (: